OpenStudy (anonymous):

Grade 10 SYSTEM OF EQUATIONS QUESTION DONT ANSWER IT, JUST HELP ME PLEASE Fran and Winston have a combined income of 80 000 dollars. One quarter of Winston's income is the same as one-sixth of Fran's income. How much does each person earn?

OpenStudy (anonymous):

what you need to do is create 2 equations involving the 2 things you want to know. in this case, it's winston's income and fran's income. can you do that?

OpenStudy (anonymous):

let x be and let y be statement? so, let x be frans income and let y be winston's income

OpenStudy (anonymous):

okay, now look at your info and see if you can come up with 2 equations involving x and y

OpenStudy (anonymous):

x+y=80 000?

OpenStudy (anonymous):

great and now another

OpenStudy (anonymous):

3/12 + 2/12 = 80 000?

OpenStudy (anonymous):

no, that's not true

OpenStudy (anonymous):

i'm having trouble making a second one

OpenStudy (anonymous):

\[\left\{f+w=80000,\frac{w}{4}=\frac{f}{6}\right\} \]

OpenStudy (anonymous):

look at this statement and see what you come up with: One quarter of Winston's income is the same as one-sixth of Fran's income.

OpenStudy (anonymous):

i just changed the fractions so they would have the same denominator...was that wrong?

OpenStudy (anonymous):

no, that wasn't wrong but how is 80,000 involved? you've already captured that bit of info in your first equation

OpenStudy (anonymous):

so it would equal 0?

OpenStudy (anonymous):

3/12 +2/12

OpenStudy (anonymous):

no, what is equivalent in the statement?

OpenStudy (anonymous):

3/12 of what? 2/12 of what?

OpenStudy (anonymous):

80 000 dollars, i think

OpenStudy (anonymous):

That will give you your second equation.

OpenStudy (anonymous):

read the statement again please

OpenStudy (anonymous):

@tomhue their income... I don't know.

OpenStudy (anonymous):

Not of 80 000 dollars, but of... yes!

OpenStudy (anonymous):

yes, their individual incomes, not the sum of them.

OpenStudy (anonymous):

They're asking what each person earned...

OpenStudy (anonymous):

right but what they're saying in the statement is that one quarter of winston's income is the same as on sixth of fran's income. what represents winston's income? what represents fran's income?

OpenStudy (anonymous):

x(2/12) + y(3/12) ?

OpenStudy (anonymous):

not their sum

OpenStudy (anonymous):

x=y

OpenStudy (anonymous):

You're not adding them. Remember, that fraction of x is the SAME as that fraction of y.

OpenStudy (anonymous):

x/6 = y/4

OpenStudy (anonymous):

why is the variable where the numerator is supposed to be?

OpenStudy (anonymous):

It's the same as saying (1/6)*(x/1) = (1/4)*(y/1). You can write it as x/6=y/4 or as (1/6)x=(1/4)y, whichever you prefer.

OpenStudy (anonymous):

like this...\[\frac{1}{4}y=\frac{1}{6}x \Rightarrow \frac{y}{4}=\frac{x}{6}\]

OpenStudy (anonymous):

pgpilot could you give tomhue a medal because he is helping me too, i appreciate the help

OpenStudy (anonymous):

OHHHHH!

OpenStudy (anonymous):

i get it now. so my two equations would be:

OpenStudy (anonymous):

1. x+y=80 000 2. y/4=x/6 and then im going to use substitution, so, 1. x=80 000-y so then, 2. y/4=80 000-y/6 ? but then how would you solve that?

OpenStudy (anonymous):

it may be easier to solve the other equation and substitute in the first.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

\[\frac{x}{6}=\frac{y}{4} \Rightarrow x = \frac{3}{2}y\]

OpenStudy (anonymous):

Correct. Just remember that you made: x = Fran's income y = Winston's income You can solve this as a proportion if you like. \[\frac{y}{4}=\frac{80 000-y}{6}\] You can use the cross products method. (top left * bottom right = top right * bottom left)

OpenStudy (anonymous):

then the first becomes \[\frac{3}{2}y+y=80,000\]

OpenStudy (anonymous):

where did the 3 over 2 come from?

OpenStudy (anonymous):

6/4 = 3/2

OpenStudy (anonymous):

x/6=y/4 (x/6)*6=(y/4)*6 x = 6y/4 x = (6/4)y x = (3/2)y

OpenStudy (anonymous):

so, @tomhue 6y+320000-4y

OpenStudy (anonymous):

6y = 320 000 -4y =, not + there. Now you can use algebra to solve this way, or substituting the reverse way like pgpilot326 is using.

OpenStudy (anonymous):

6y+4y=320 000 10y= 320 000 y=32 000 so, Winston earns 32 000

OpenStudy (anonymous):

Yup! Now, how much does Fran earn? Remember that x + y = 80000. You just found the value of y.

OpenStudy (anonymous):

so then, x+32 000=80 000 x=80 000-32 000 x= 48 0000 Fran earns 48 000

OpenStudy (anonymous):

wow! Thanks for the help!

OpenStudy (anonymous):

Correct. You got your answer. Winston 32000, Fran 48000. You're welcome.

OpenStudy (anonymous):

is there any advice you can give me on these types of questions? because I have five more like this..and I struggle with them...

OpenStudy (anonymous):

Turn the names of people/things into variables. An amount "OF" a person's income/whatever means you multiply that amount by the variable. 1/6 of Fran's income means 1/6 * x. There will usually be one equation adding the two variables or multiples of them to make a number. The other will probably (but not definitely) have variables on both sides.

OpenStudy (anonymous):

could we try one more? do you have time?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

thanks so much. I'll just type it

OpenStudy (anonymous):

In one city, taxi company A charges $5 plus $0.35/km travelled. Taxi company B charges $3.50 plus $0.50/km. For what distance is the charge the same using either taxi company? In what situations would you choose company A?

OpenStudy (anonymous):

there isn't a total distance...and how do I find out the kilometers?

OpenStudy (anonymous):

Turn kilometer into a variable. "Of" and "per" and "/" all are indicators for multiplication when used with a number and a unit/thing/person.

OpenStudy (anonymous):

Let "x" be km but what would "y" be?

OpenStudy (anonymous):

i know the money...

OpenStudy (anonymous):

Actually, this doesn't even look like it's a system. You know the base cost of each company, plus the per km cost. A = base + additional*x B = base + additional*x You're looking for the scenario when A = B.

OpenStudy (anonymous):

so...its not a system of equations question?

OpenStudy (anonymous):

You could call it a system if you wanted, but it's only one-variable. It's still worth knowing how to do.

OpenStudy (anonymous):

A= 5+.35x B= 3.50+.50x

OpenStudy (anonymous):

so how do I find x?

OpenStudy (anonymous):

Remember, you want A = B. You can use substitution.

OpenStudy (anonymous):

one second, brb

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

5+.35x=3.50+.50x

OpenStudy (anonymous):

Yup! Then you solve for x like you normally would.

OpenStudy (anonymous):

5-3.50=.50x-.35x

OpenStudy (anonymous):

1.5=.15x

OpenStudy (anonymous):

10 x.

OpenStudy (anonymous):

THANKS AGAIN!

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

so now i have to find the same distances

OpenStudy (anonymous):

a= 5+.35 (10) a= 5+3.5 a=8.5 b=3.50+.50(10) b=3.50+ 5 b=8.5 oh wait it was 10...

OpenStudy (anonymous):

Oh, yeah. X was the number of kilometers. It was a one-variable system. I was just using A and B to show the cost of the companies. You wanted the same cost, so A=B.

OpenStudy (anonymous):

i'm going to check for nine to see which is cheaper. a=5+.35(9) a=5+3.15 a=8.15 b=3.50+.50(9) b=3.50+4.5 b=8 so, I would choose b because its cheaper before 10 km a=5+.35(11) a=5+3.85 a= 8.85 b= 3.50+.50(11) b=3.50+5.5 b=9 SO I WOULD CHOOSE A IF I NEED TO GO 10 KM OR MORE. because its cheaper after 10km.

OpenStudy (anonymous):

Yup! If you wanted to, you could just take the slope of the equation (coefficient of the variable in these cases). The equation with the smaller slope is cheaper in the long run. Try not to be fooled into thinking that these problems are harder than they actually are. If it's really unbearable, just save the confusing word problems for the end of the test. Go back to them.

OpenStudy (anonymous):

Thank you so much. You have helped me alot.